OpenStudy (anonymous):

Find the X and Y intercepts m=2 (-3,4)

Nnesha (nnesha):

y=mx+b slope form equation replace m by 2 and then replace x and y by their values to find y -intercept

OpenStudy (campbell_st):

this is a 2 part problem find the equation of the line the slope intercept form of a line is y=mx + b you have m = 2 and using the point x = -3 and y =4 you can substitute the values to find b, and the equation of the line.

OpenStudy (anonymous):

sooooo 4 = 2 x - 3 + b?

Nnesha (nnesha):

replace x by -3

Nnesha (nnesha):

yeah that's right

OpenStudy (anonymous):

i meant 2 times -3

Nnesha (nnesha):

yeah

Nnesha (nnesha):

solve for b

OpenStudy (anonymous):

10 = b

Nnesha (nnesha):

perfect so 10 is what ??

OpenStudy (anonymous):

the y intercept which is 0,10

Nnesha (nnesha):

yes right now you have to find x -intercept x -intercept where line cross x'axis where y is 0 so replace y by 0 and then solve for x

Nnesha (nnesha):

wait do you know how to write equation right y= mx+b first replace m by 2 and b by 10

OpenStudy (anonymous):

0 = 2x+10?

Nnesha (nnesha):

yesh!! now solve for x

OpenStudy (anonymous):

-5 = x which is -5,0

Nnesha (nnesha):

yesh \(\huge\color{green}{{\rm good}~\rm :)job!!}\)

OpenStudy (anonymous):

ty :D

Nnesha (nnesha):

my pleasure :)